FTMW AND FTMW-MMW DOUBLE RESONANCE SPECTROSCOPY OF AN OPEN-SHELL RADICAL COMPLEX, AR-HO_{2}.

\underline{KOHSUKE SUMA}, WATARU FUNATO, YOSHIHIRO SUMIYOSHI, AND YASUKI ENDO, Department of Basic Sciences, Graduate School of Arts and Sciences, The University of Tokyo, Komaba, Meguro-ku, Tokyo 153-8902, Japan.

Although a number of molecular complexes have been studied extensively for a long time, studies on open-shell radical complexes are relatively scarce due to its difficulty of production. Besides, most of them deal with relatively simple rare-gas(Rg)-diatomic systems. We report the first microwave observation of the Rg-open-shell triatomic species, Ar-HO_{2}. The a-type and b-type transitions of Ar-HO_{2} have been observed using Fourier transform microwave (FTMW) and FTMW-millimeter wave double resonance spectroscopy. The r_{0} structure of Ar-HO_{2} has been determined from the rotational constants by fixing the structure of the HO_{2} monomer. The determined molecular structure is cis planar, where the argon atom is attached to the hydrogen atom of HO_{2}; r_{0}(ArH) = 2.622Å, and θ_{0}(ArHO)= 153.9^{∘}. r_{0}(ArH) of Ar-HO_{2} is shoter than that of Ar-HO, 2.776Å^{a}, indicating that the Ar-HO_{2} bond is relatively rigid. In fact, the binding energy of Ar-HO_{2}, 292cm^{−1}, calculated by RCCSD(T)/aug-cc-pVTZ + bond function is much larger than that of Ar-HO or other closed-shell Rg-triatomic species. On the other hand, the determined fine and hyperfine structures are almost identical to those of the HO_{2} monomer. The large binding energy may be ascribed to the effect of the low-lying first excited state of the HO_{2} monomer.

aY. Ohshima, M. Iida, and Y. Endo, J. Chem. Phys. 95, 7001 (1991)